# On geometric spectral functionals

### v1

@@ 3.1 The spin Dirac operator
(cf. [9, Theorem 2.2]) For the Dirac-type operator $D=D_{0}+B$ with $B$ given by Eq. (3.19), the torsion functional reads
$\mathscr{T}_{D}(u,v,w)=-\frac{3i}{2}\nu_{n-1}\dim(V)\int_{M}\mathrm{dvol}_{g}\,u_{a}v_{b}w_{c}{}^{A}T^{0}_{abc}.$ (3.23)
Proof 3.20.
It follows directly from Proposition 3.10 and by using (3.21).
Theorem 3.21.
The scalar curvature functional for the torsion perturbation of the spin Dirac operator, $B=-\frac{i}{8}{}^{A}T_{abc}$ reads
$\mathscr{R}_{D}(f)=\frac{2^{m}(n-2)\nu_{n-1}}{24}\int_{M}\mathrm{dvol}_{g}f\Bigl{[}-R+\frac{9}{4}{}^{A}T_{abc}^{0}{}^{A}T_{abc}^{0}\Bigr{]}.$ (3.24)
Proof 3.22.
This follows directly from
$\mathrm{Tr}(-12B_{0}^{2}+6\gamma^{a}\{\gamma^{a},B_{0}\}B_{0})=2^{m}\cdot\frac{9}{4}{}^{A}T_{abc}^{0}{}^{A}T_{abc}^{0},$ (3.25)
Remark 3.23.
The result presented in Theorem 3.21 is consistent with those found in the existing literature, for example, in [13, p. 879] (noting that their definition of $(T^{0})^{2}$ differs from ours by a factor of $\frac{1}{6}$), as well as in [16, Proposition 5.4] and [15, Section 5.2]. We note that in the latter works, the term “torsion” refers to what is technically the contorsion, which introduces an additional factor of $\frac{1}{4}$ compared to our formulation. However, there is an inconsistency with the formula given in [1, Lemma 3.1], where a factor related to the rank of the corresponding bundle appears, which should not be present.
3.2 The Hodge-Dirac operator
We recall here the construction of the Hodge-Dirac operator from [8]; however, we introduce a family of Dirac-type operators extended by torsion.
We review the basics of our notation from [8]. We use multi-index notation to represent differential forms and operations on them compactly. For $0\leq l\leq n$, a differential $l$-form $\omega$ is denoted by $\omega=\sum_{J}\omega_{J}dx^{J}$, where $J$ is an ordered multi-index. Raising and lowering the degrees of differential forms are performed using the operators $\lambda^{p}_{+}$ and $\lambda_{-}^{p}$, respectively, whose components are given by
$(\lambda^{p}_{+})^{I}_{J}=\varepsilon^{I}_{pJ},\qquad(\lambda^{p}_{-})^{I}_{J}=\varepsilon^{pI}_{J}.$ (3.26)
They act on basis forms $e^{J}$ as
$\lambda^{p}_{+}e^{J}=\varepsilon^{I}_{pJ}e^{I},\qquad\lambda^{p}_{-}e^{J}=\varepsilon^{pI}_{J}e^{I},$ (3.27)
where
$\varepsilon^{I}_{pJ}=\begin{cases}(-1)^{\mathrm{sgn}(\pi)},\quad\exists\mbox{ permutation }\pi\mbox{ s.t. }pJ=\pi(I),\\ 0,\qquad\mbox{otherwise}.\end{cases}$ (3.28)
They satisfy
$\{\lambda_{\pm}^{p},\lambda_{\pm}^{r}\}=0,\qquad\{\lambda^{p}_{+},\lambda^{r}_{-}\}=\delta_{pr}\mathrm{id},$ (3.29)
and the operators $\gamma^{p}=-i(\lambda^{p}_{+}-\lambda^{p}_{-})$ generates Clifford algebra $\{\gamma^{p},\gamma^{r}\}=\delta_{pr}\mathrm{id}$.
The Hodge-Dirac operator, $D_{0}=d+d^{\ast}$, was studied from the spectral perspective in [8]. Here we study the perturbed Hodge-Dirac operator, $D=D_{0}+B$ with
$B$ $=\frac{1}{2}T_{ijk}\left(\lambda_{+}^{j}\lambda_{+}^{i}\lambda_{-}^{k}+\lambda_{+}^{k}\lambda_{-}^{i}\lambda_{-}^{j}\right)$ (3.30)
$=\frac{1}{2}T_{ijk}\Bigl{[}\delta_{ik}(\lambda^{j}_{+}+\lambda^{j}_{-})-\lambda^{j}_{+}\lambda^{k}_{-}\lambda^{i}_{+}+\lambda^{j}_{-}\lambda^{k}_{+}\lambda^{i}_{-}\Bigr{]}.$
First, we demonstrate that the above $B$ can be considered as a generalization of the Hodge-Dirac operator with torsion. Since the exterior derivative $d$ is expressed using the Levi-Civita connection $d=\sum_{i}dx^{i}\wedge\nabla^{\mathrm{LC}}_{i}$, we construct a first-order differential operator $\tilde{d}=\sum_{i}dx^{i}\wedge\nabla_{i}$ for an arbitrary metric compatible linear connection $\nabla$. Explicitly, on a one-form $u$,
$(\tilde{d}u)_{J}=\left(\partial_{j}u_{I}-\Gamma^{k}_{jk}u_{I}+\varepsilon^{iK}_{kI}\Gamma^{k}_{ji}u_{K}\right)\varepsilon^{jI}_{J}=\left(\partial_{j}u_{I}-\varepsilon^{iA}_{I}\varepsilon^{K}_{kA}\Gamma^{k}_{ji}u_{K}\right)\varepsilon^{jI}_{J},$ (3.31)
$\tilde{d}u=\lambda_{+}^{j}\partial_{j}u+\frac{1}{2}\lambda_{+}^{j}\lambda_{+}^{i}\lambda_{-}^{k}T_{ijk}u.$ (3.32)
The Hodge-Dirac operator with torsion is then defined as $D=\tilde{d}+\tilde{d}^{\ast}$, which for $T_{ijk}=0$ is the usual Hodge-Dirac $d+d^{*}$. Using the definition of an adjoint operator $A^{\ast}$ of an operator $A$, in normal coordinates we get
$(A^{\ast})_{LJ}=\overline{A_{JL}}+o({\bf x}).$ (3.33)
Therefore,
$\tilde{d}^{\ast}=d^{\ast}+\frac{1}{2}\lambda_{+}^{k}\lambda_{-}^{i}\lambda_{-}^{j}T_{ijk}.$ (3.34)
Observe that the operator $D$ is a Dirac-type operator as defined in Definition 2.4; thus, it yields the standard metric functional on $M$ (up to a multiplicative constant). We also note that a Hodge-Dirac operator coupled to antisymmetric torsion appears in [17]; however, comparing it to the construction presented here is not straightforward.
Proposition 3.24.
For the Hodge-Dirac operator with torsion (3.30), the Einstein functional densities at point $x$ in normal coordinates reads,
$\mathscr{G}_{D}(u,w)(x)$ $=\mathscr{G}_{D_{0}}(u,w)(x)+3\nu_{n-1}2^{n-3}u_{a}\Biggl{[}-4w_{bc}{}^{A}T^{0}_{abc}-2w_{b}{}^{A}T^{c}_{abc}$ (3.35)
$-\frac{4}{3}w_{a}T^{0}_{jii}T^{0}_{jkk}+w_{b}{}^{A}T^{0}_{cjk}{}^{A}T^{0}_{djk}\Bigl{(}\frac{1}{2}\delta_{ab}\delta_{cd}-3\delta_{ac}\delta_{bd}\Bigr{)}\Biggr{]}.$
where ${}^{A}T_{ijk}=\frac{1}{3}(T_{ijk}+T_{kij}+T_{jki})$ and $\mathscr{G}_{D_{0}}(u,w)$ is given by [9, Proposition 3.3].
@@ 3.2 The Hodge-Dirac operator
The following identities hold
$\mathrm{Tr}(\gamma^{a}\lambda^{i}_{\pm})=\pm i2^{n-1}\delta_{ai},$
$\mathrm{Tr}(\gamma^{a}\lambda^{j}_{\mp}\lambda^{k}_{\pm}\lambda^{i}_{\mp})=\mp i2^{n-2}(\delta_{ai}\delta_{kj}+\delta_{aj}\delta_{ki}),$
$\mathrm{Tr}(\gamma^{a}\gamma^{b}\gamma^{c}\lambda^{i}_{\pm})=\pm i2^{n-1}(\delta_{ai}\delta_{bc}-\delta_{bi}\delta_{ac}+\delta_{ci}\delta_{ab}),$
$\begin{split}\mathrm{Tr}(\gamma^{a}\gamma^{b}\gamma^{c}\lambda^{j}_{\pm}\lambda^{k}_{\mp}\lambda^{i}_{\pm})=\pm i&2^{n-3}\Bigl{[}2\delta_{bc}(\delta_{ai}\delta_{kj}+\delta_{aj}\delta_{ki})-2\delta_{ac}(\delta_{bi}\delta_{kj}+\delta_{bj}\delta_{ki})\\ &+2\delta_{ab}(\delta_{ic}\delta_{kj}+\delta_{cj}\delta_{ki})+\delta_{ka}\varepsilon^{ji}_{bc}+\delta_{kb}\varepsilon^{ji}_{ca}+\delta_{kc}\varepsilon^{ji}_{ab}\Bigr{]},\end{split}$
$\mathrm{Tr}(\gamma^{a}\gamma^{b}\lambda^{j}_{\mp}\lambda^{i}_{\mp})=-2^{n-2}\varepsilon^{ij}_{ab},$
$\mathrm{Tr}(\gamma^{a}\gamma^{b}\gamma^{c}\gamma^{d}\lambda^{j}_{\pm}\lambda^{i}_{\pm})=2^{n-2}\Bigl{(}\delta_{ab}\varepsilon^{ji}_{cd}+\delta_{ac}\varepsilon^{ji}_{db}+\delta_{ad}\varepsilon^{ji}_{bc}+\delta_{bc}\varepsilon^{ji}_{ad}+\delta_{bd}\varepsilon^{ji}_{ca}+\delta_{cd}\varepsilon^{ji}_{ab}\Bigr{)}.$
Furthermore, by straightforward computations, we have
@@
With $B$ given by (3.30) the following identities hold
$\mathrm{Tr}(\gamma^{a}B)=0,$
$\mathrm{Tr}(\gamma^{a}\gamma^{b}\gamma^{c}B)=i2^{n-2}(T_{abc}+T_{cab}+T_{bca}),$
$\mathrm{Tr}(\gamma^{a}\gamma^{b}\{\gamma^{c},B\})=i2^{n-1}(T_{abc}+T_{cab}+T_{bca}),$
$\mathrm{Tr}(\gamma^{a}\gamma^{b}\gamma^{c}\gamma^{d}\{\gamma^{d},B\})=3i\cdot 2^{n-1}(T_{abc}+T_{cab}+T_{bca}),$
$\mathrm{Tr}(B^{2})=2^{n-2}T_{baa}T_{bcc}+2^{n-3}T_{abc}T_{abc},$
$\mathrm{Tr}(\gamma^{a}\{\gamma^{b},B\}B)=-2^{n-2}T_{ajk}T_{bkj}+2^{n-2}T_{ajk}T_{bjk}+2^{n-3}T_{jka}T_{jkb},$
$\mathrm{Tr}(\gamma^{c}\gamma^{a}\{\gamma^{b},B\}\{\gamma^{c},B\})=2^{n-1}[T_{jkb}(T_{jka}+T_{ajk})+T_{bkj}(T_{kja}+2T_{akj}-2T_{ajk})].$
Proof 3.27 (Proof of Proposition 3.24).
It follows from combining Proposition 3.6 and Lemma 3.26.
@@
For the Hodge-Dirac operator with torsion (3.30), the torsion functional reads
$\mathscr{T}_{D}(u,v,w)=-3i\nu_{n-1}2^{n-1}\int_{M}\mathrm{dvol}_{g}u_{a}v_{b}w_{c}({}^{A}T^{0}_{abc}).$ (3.36)
Proof 3.29.
It is an immediate consequence of Lemma 3.26.
@@
For the Hodge-Dirac operator with torsion (3.30), the scalar curvature functional reads
$\begin{split}\mathscr{R}_{D}(f)&=\frac{2^{n-3}}{3}(n-2)\nu_{n-1}\int_{M}\mathrm{dvol}_{g}f\Bigl{[}-R-3T^{0}_{baa}T^{0}_{bcc}+\frac{3}{4}\left(T^{0}_{abc}T^{0}_{abc}+2T^{0}_{abc}T^{0}_{cab}\right)\Bigr{]}\\ &=\frac{2^{n-3}}{3}(n-2)\nu_{n-1}\int_{M}\mathrm{dvol}_{g}f\Bigl{[}-R-3T^{0}_{baa}T^{0}_{bcc}+\frac{9}{4}{}^{A}T^{0}_{abc}{}^{A}T^{0}_{abc}\Bigr{]}.\end{split}$ (3.37)
Proof 3.32.
From [8, Lemma 2.1], we immediately infer that
$\mathrm{Tr}(P^{0}_{aa})=\frac{2^{n}}{3}R,\qquad\mathrm{Tr}(Q^{0})=\frac{2^{n-2}}{3}R.$ (3.38)
Combining this with Lemma 3.26 and Proposition 3.13, the result follows.
@@
We observe that if the vector part of the torsion vanishes, the scalar curvature functional for the Hodge-Dirac operator in Proposition 3.31 aligns with the scalar curvature functional for the spin Dirac operator from Theorem 3.21, up to an overall coefficient. We can summarize the result for the effective density of the functional scalar curvature in the presence of torsion, $R_{T}$, as
$R_{T}=R-\frac{9}{4}T^{A0}_{abc}T^{A0}_{abc}+\begin{cases}0,\qquad\mbox{spin Dirac},\\ 3T^{0}_{baa}T^{0}_{bcc},\quad\mbox{Hodge-Dirac}.\end{cases}$ (3.39)
Note that for the vector torsion, which is permitted in the case of the Hodge-Dirac operator, the sign differs from that of the antisymmetric part of the torsion, and thus the functional may not have extrema.

### v2

@@ 3.1 The spin Dirac operator
Proposition 3.11 (cf. [8, Theorem 2.2]).
For the Dirac-type operator $D=D_{0}+B$ with $B$ given by equation (3.2), the torsion functional reads
$\mathscr{T}_{D}(u,v,w)=-\frac{3{\rm i}}{2}\nu_{n-1}\dim(V)\int_{M}\mathrm{dvol}_{g}u_{a}v_{b}w_{c}{}^{A}T^{0}_{abc}.$
Proof.
It follows directly from Proposition 3.6 and by using (3.3). ∎
Theorem 3.12.
The scalar curvature functional for the torsion perturbation of the spin Dirac operator, $B=-\frac{\rm i}{8}{}^{A}T_{abc}$ reads
$\mathscr{R}_{D}(f)=\frac{2^{m}(n-2)\nu_{n-1}}{24}\int_{M}\mathrm{dvol}_{g}f\left[-R+\frac{9}{4}{}^{A}T_{abc}^{0}{}^{A}T_{abc}^{0}\right].$
Proof.
This follows directly from
$\mathrm{Tr}\bigl(-12B_{0}^{2}+6\gamma^{a}\{\gamma^{a},B_{0}\}B_{0}\bigr)=2^{m}\cdot\frac{9}{4}{}^{A}T_{abc}^{0}{}^{A}T_{abc}^{0}.$
Remark 3.13.
The result presented in Theorem 3.12 is consistent with those found in the existing literature, for example, in [13, p. 879] (noting that their definition of $\bigl(T^{0}\bigr)^{2}$ differs from ours by a factor of $\frac{1}{6}$), as well as in [17, Proposition 5.4] and [16, Section 5.2]. We note that in the latter works, the term “torsion” refers to what is technically the contorsion, which introduces an additional factor of $\frac{1}{4}$ compared to our formulation. However, there is an inconsistency with the formula given in [1, Lemma 3.1], where a factor related to the rank of the corresponding bundle appears, which should not be present.
3.2 The Hodge–Dirac operator
We recall here the construction of the Hodge–Dirac operator from [9]; however, we introduce a family of Dirac-type operators extended by torsion.
We review the basics of our notation from [9]. We use multi-index notation to represent differential forms and operations on them compactly. For $0\leq l\leq n$, a differential $l$-form $\omega$ is denoted by $\omega=\sum_{J}\omega_{J}{\rm d}x^{J}$, where $J$ is an ordered multi-index. Raising and lowering the degrees of differential forms are performed using the operators $\lambda^{p}_{+}$ and $\lambda_{-}^{p}$, respectively, whose components are given by
$\bigl(\lambda^{p}_{+}\bigr)^{I}_{J}=\varepsilon^{I}_{pJ},\qquad\bigl(\lambda^{p}_{-}\bigr)^{I}_{J}=\varepsilon^{pI}_{J}.$
They act on basis forms $e^{J}$ as $\lambda^{p}_{+}e^{J}=\varepsilon^{I}_{pJ}e^{I}$, $\lambda^{p}_{-}e^{J}=\varepsilon^{pI}_{J}e^{I}$, where
$\varepsilon^{I}_{pJ}=\begin{cases}(-1)^{\mathrm{sgn}(\pi)},&\exists\mbox{ permutation }\pi\mbox{ s.t. }pJ=\pi(I),\\ 0,&\mbox{otherwise}.\end{cases}$
They satisfy $\bigl\{\lambda_{\pm}^{p},\lambda_{\pm}^{r}\bigr\}=0$, $\bigl\{\lambda^{p}_{+},\lambda^{r}_{-}\bigr\}=\delta_{pr}\mathrm{id}$, and the operators $\gamma^{p}=-i\bigl(\lambda^{p}_{+}-\lambda^{p}_{-}\bigr)$ generates Clifford algebra $\{\gamma^{p},\gamma^{r}\}=\delta_{pr}\mathrm{id}$.
The Hodge–Dirac operator, $D_{0}=d+d^{\ast}$, was studied from the spectral perspective in [9]. Here we study the perturbed Hodge–Dirac operator, $D=D_{0}+B$ with
$B$ $=\frac{1}{2}T_{ijk}\bigl(\lambda_{+}^{j}\lambda_{+}^{i}\lambda_{-}^{k}+\lambda_{+}^{k}\lambda_{-}^{i}\lambda_{-}^{j}\bigr)=\frac{1}{2}T_{ijk}\bigl[\delta_{ik}\bigl(\lambda^{j}_{+}+\lambda^{j}_{-}\bigr)-\lambda^{j}_{+}\lambda^{k}_{-}\lambda^{i}_{+}+\lambda^{j}_{-}\lambda^{k}_{+}\lambda^{i}_{-}\bigr].$ (3.4)
First, we demonstrate that the above $B$ can be considered as a generalization of the Hodge–Dirac operator with torsion. Since the exterior derivative $d$ is expressed using the Levi-Civita connection ${\rm d}=\sum_{i}{\rm d}x^{i}\wedge\nabla^{\mathrm{LC}}_{i}$, we construct a first-order differential operator $\tilde{\rm d}=\sum_{i}{\rm d}x^{i}\wedge\nabla_{i}$ for an arbitrary metric compatible linear connection $\nabla$. Explicitly, on a one-form $u$,
$\bigl(\tilde{\rm d}u\bigr)_{J}=\bigl(\partial_{j}u_{I}-\Gamma^{k}_{jk}u_{I}+\varepsilon^{iK}_{kI}\Gamma^{k}_{ji}u_{K}\bigr)\varepsilon^{jI}_{J}=\bigl(\partial_{j}u_{I}-\varepsilon^{iA}_{I}\varepsilon^{K}_{kA}\Gamma^{k}_{ji}u_{K}\bigr)\varepsilon^{jI}_{J},$
$\tilde{\rm d}u=\lambda_{+}^{j}\partial_{j}u+\frac{1}{2}\lambda_{+}^{j}\lambda_{+}^{i}\lambda_{-}^{k}T_{ijk}u.$
The Hodge–Dirac operator with torsion is then defined as $D=\tilde{\rm d}+\tilde{\rm d}^{\ast}$, which for $T_{ijk}=0$ is the usual Hodge–Dirac $d+d^{*}$. Using the definition of an adjoint operator $A^{\ast}$ of an operator $A$, in normal coordinates we get $(A^{\ast})_{LJ}=\overline{A_{JL}}+o({\bf x})$. Therefore,
$\tilde{\rm d}^{\ast}=d^{\ast}+\frac{1}{2}\lambda_{+}^{k}\lambda_{-}^{i}\lambda_{-}^{j}T_{ijk}.$
Observe that the operator $D$ is a Dirac-type operator as defined in Definition 2.4; thus, it yields the standard metric functional on $M$ (up to a multiplicative constant). We also note that a Hodge–Dirac operator coupled to antisymmetric torsion appears in [18]; however, comparing it to the construction presented here is not straightforward.
@@ 3.2 The Hodge–Dirac operator
For the Hodge–Dirac operator with torsion (3.4), the Einstein functional densities at point $x$ in normal coordinates reads
$\mathscr{G}_{D}(u,w)(x)={}$ $\mathscr{G}_{D_{0}}(u,w)(x)+3\nu_{n-1}2^{n-3}u_{a}\left[-4w_{bc}{}^{A}T^{0}_{abc}-2w_{b}{}^{A}T^{c}_{abc}\right$.
$\left.-\frac{4}{3}w_{a}T^{0}_{jii}T^{0}_{jkk}+w_{b}{}^{A}T^{0}_{cjk}{}^{A}T^{0}_{djk}\left(\frac{1}{2}\delta_{ab}\delta_{cd}-3\delta_{ac}\delta_{bd}\right)\right],$
where $^{A}T_{ijk}=\frac{1}{3}(T_{ijk}+T_{kij}+T_{jki})$ and $\mathscr{G}_{D_{0}}(u,w)$ is given by [8, Proposition 3.3].
@@
The following identities hold:
$\mathrm{Tr}\bigl(\gamma^{a}\lambda^{i}_{\pm}\bigr)=\pm{\rm i}2^{n-1}\delta_{ai},\qquad\mathrm{Tr}\bigl(\gamma^{a}\lambda^{j}_{\mp}\lambda^{k}_{\pm}\lambda^{i}_{\mp}\bigr)=\mp{\rm i}2^{n-2}(\delta_{ai}\delta_{kj}+\delta_{aj}\delta_{ki}),$
$\mathrm{Tr}\bigl(\gamma^{a}\gamma^{b}\gamma^{c}\lambda^{i}_{\pm}\bigr)=\pm{\rm i}2^{n-1}(\delta_{ai}\delta_{bc}-\delta_{bi}\delta_{ac}+\delta_{ci}\delta_{ab}),$
$\mathrm{Tr}\bigl(\gamma^{a}\gamma^{b}\gamma^{c}\lambda^{j}_{\pm}\lambda^{k}_{\mp}\lambda^{i}_{\pm}\bigr)=\pm 2^{n-3}\bigl[2\delta_{bc}(\delta_{ai}\delta_{kj}+\delta_{aj}\delta_{ki})-2\delta_{ac}(\delta_{bi}\delta_{kj}+\delta_{bj}\delta_{ki})$
[…]
$\mathrm{Tr}\bigl(\gamma^{a}\gamma^{b}\lambda^{j}_{\mp}\lambda^{i}_{\mp}\bigr)=-2^{n-2}\varepsilon^{ij}_{ab},$
$\mathrm{Tr}\bigl(\gamma^{a}\gamma^{b}\gamma^{c}\gamma^{d}\lambda^{j}_{\pm}\lambda^{i}_{\pm}\bigr)=2^{n-2}\bigl(\delta_{ab}\varepsilon^{ji}_{cd}+\delta_{ac}\varepsilon^{ji}_{db}+\delta_{ad}\varepsilon^{ji}_{bc}+\delta_{bc}\varepsilon^{ji}_{ad}+\delta_{bd}\varepsilon^{ji}_{ca}+\delta_{cd}\varepsilon^{ji}_{ab}\bigr).$
Furthermore, by straightforward computations, we have the following.
@@
With $B$ given by (3.4) the following identities hold:
$\mathrm{Tr}(\gamma^{a}B)=0,\qquad\mathrm{Tr}\bigl(\gamma^{a}\gamma^{b}\gamma^{c}B\bigr)={\rm i}2^{n-2}(T_{abc}+T_{cab}+T_{bca}),$
$\mathrm{Tr}\bigl(\gamma^{a}\gamma^{b}\{\gamma^{c},B\}\bigr)={\rm i}2^{n-1}(T_{abc}+T_{cab}+T_{bca}),$
$\mathrm{Tr}\bigl(\gamma^{a}\gamma^{b}\gamma^{c}\gamma^{d}\bigl\{\gamma^{d},B\bigr\}\bigr)=3{\rm i}\cdot 2^{n-1}(T_{abc}+T_{cab}+T_{bca}),$
$\mathrm{Tr}\bigl(B^{2}\bigr)=2^{n-2}T_{baa}T_{bcc}+2^{n-3}T_{abc}T_{abc},$
$\mathrm{Tr}\bigl(\gamma^{a}\bigl\{\gamma^{b},B\bigr\}B\bigr)=-2^{n-2}T_{ajk}T_{bkj}+2^{n-2}T_{ajk}T_{bjk}+2^{n-3}T_{jka}T_{jkb},$
$\mathrm{Tr}\bigl(\gamma^{c}\gamma^{a}\bigl\{\gamma^{b},B\bigr\}\{\gamma^{c},B\}\bigr)=2^{n-1}[T_{jkb}(T_{jka}+T_{ajk})+T_{bkj}(T_{kja}+2T_{akj}-2T_{ajk})].$
Proof of Proposition 3.14.
It follows from combining Proposition 3.4 and Lemma 3.16. ∎
@@
For the Hodge–Dirac operator with torsion (3.4), the torsion functional reads
$\mathscr{T}_{D}(u,v,w)=-3{\rm i}\nu_{n-1}2^{n-1}\int_{M}\mathrm{dvol}_{g}u_{a}v_{b}w_{c}\bigl({}^{A}T^{0}_{abc}\bigr).$
Proof.
It is an immediate consequence of Lemma 3.16. ∎
@@
For the Hodge–Dirac operator with torsion (3.4), the scalar curvature functional reads
$\mathscr{R}_{D}(f)$ $=\frac{2^{n-3}}{3}(n-2)\nu_{n-1}\int_{M}\mathrm{dvol}_{g}f\left[-R-3T^{0}_{baa}T^{0}_{bcc}+\frac{3}{4}\bigl(T^{0}_{abc}T^{0}_{abc}+2T^{0}_{abc}T^{0}_{cab}\bigr)\right]$
$=\frac{2^{n-3}}{3}(n-2)\nu_{n-1}\int_{M}\mathrm{dvol}_{g}f\left[-R-3T^{0}_{baa}T^{0}_{bcc}+\frac{9}{4}{}^{A}T^{0}_{abc}{}^{A}T^{0}_{abc}\right].$
Proof.
From [9, Lemma 2.1], we immediately infer that
$\mathrm{Tr}\bigl(P^{0}_{aa}\bigr)=\frac{2^{n}}{3}R,\qquad\mathrm{Tr}\bigl(Q^{0}\bigr)=\frac{2^{n-2}}{3}R.$
Combining this with Lemma 3.16 and Proposition 3.8, the result follows. ∎
@@
We observe that if the vector part of the torsion vanishes, the scalar curvature functional for the Hodge–Dirac operator in Proposition 3.19 aligns with the scalar curvature functional for the spin Dirac operator from Theorem 3.12, up to an overall coefficient. We can summarize the result for the effective density of the functional scalar curvature in the presence of torsion, $R_{T}$, as
$R_{T}=R-\frac{9}{4}T^{A0}_{abc}T^{A0}_{abc}+\begin{cases}0,&\mbox{spin Dirac},\\ 3T^{0}_{baa}T^{0}_{bcc},&\mbox{Hodge--Dirac}.\end{cases}$
Note that for the vector torsion, which is permitted in the case of the Hodge–Dirac operator, the sign differs from that of the antisymmetric part of the torsion, and thus the functional may not have extrema.